OpenStudy (anonymous):

what would be formula of this graphic (y=x^2-1) if you move 2 units to the left and 3 units up. can i have some help i think is ti y+2=x^2+2 am i right or wrong please help.

myininaya (myininaya):

y=a(x-h)^2+k is the vertex form of a parabola if h>0 we move it right h units if h<0 we move it left h units if k>0 we move it up k units if k<0 we move it down k units Example y=(x+3)^2+4 is the graph of y=x^2 moved left 3 units and up 4 units Example y=(x+2)^2-4 is the graph of y=x^2 moved left 2 units and down 4 units Example y=(x-2)^2 is the graph of y=x^2 moved right 2 units

myininaya (myininaya):

So if you are moving it from left to right you are changing the inside of that square thing so can you tell me what the inside of that square thing will look like if we move from 2 units the left?

OpenStudy (anonymous):

yeah i would be x^2+2-1 = x^2+1 am i right?

myininaya (myininaya):

no.. the thing inside the square should change since we are moving left/right

myininaya (myininaya):

we will deal with outside when you can figure out the inside